OpenStudy (anonymous):

How do I solve x+2y=3 & x-y=6 using elimination?

OpenStudy (anonymous):

{ x + 2y = 3 { x - y = 6 Well since y could already be in subtraction/addition form I would multiply the second equation by a negative to. Remember, when using elimination, you're trying to use one variable. SO now, try it this way? 2 {x - y = 6} AFTER, your systems should look like { x + 2y = 3 { 2x - 2y = 12 Can you do it from there?

OpenStudy (anonymous):

So, I should try to solve one of the equations & then substitute that number into both equations?

OpenStudy (anonymous):

I think I see...you make the equations so that either x or y will cancel out when you subtract?

OpenStudy (anonymous):

Noo, when you're using elimination you don't substitute in anything. You see how on top of eachother there's a positive 2y and a negative 2y? What do they do? Cancel eachother out, right? In elimination you have to either subtract or add the top/bottom. Like this.

OpenStudy (anonymous):

Ok. I understand now. Thank you!